# FCMR: Robust Evaluation of Financial Cross-Modal Multi-Hop Reasoning

### v1

@@ 7 Conclusion
We introduce FCMR, a new benchmark designed to evaluate the cross-modal multi-hop reasoning ability of MLLMs. We evaluate the performance of various state-of-the-art MLLMs, revealing that current models continue to struggle with reasoning across different modalities. As future work, we plan to develop methods to enhance the performance of MLLMs based on the observations and analyses presented in this study.
@@
Heavy Reliance on Manual Analysis
Our analysis required extensive manual sampling and verification to ensure high-quality insights. Future work could aim to automate this process, enhancing efficiency and scalability.
@@
Room for Prompt Optimization
Providing optimized prompts for each MLLM could improve performance in evaluation. However, our top priority in this work is to test all models under equal conditions. Additionally, recent models are known to be optimized with diverse prompts during the post-training phase, making them increasingly robust to variations in instructions and prompts. For example, our experiments show that questions requiring complex reasoning often naturally elicit strategies similar to CoT, even when such strategies are not explicitly mentioned.
@@ Appendix A Details of Material Cross-Modal Multi-Hop Reasoning (Material-CMR)
The proposed CMRGen framework is easily adaptable to various domains, and as an example, we applied it to the field of Material Science to create the Material Cross-Modal Multi-Hop Reasoning (Material-CMR) dataset. We construct a Table Source containing material properties and a Text Source describing the materials through The Materials Project. The Materials Projecthttps://next-gen.materialsproject.org/ is an initiative that accelerates materials discovery by providing researchers with computational data and tools to predict material properties, enabling more targeted and efficient experimental research.
@@ B.1 Table Source
Considering the practicality of cross-modal multi-hop reasoning, we utilize annual financial statements, an essential element in real corporate analysis, as the table data source. Wharton Research Data Services (WRDS) Compustathttps://wrds-www.wharton.upenn.edu/pages/grid-items/compustat-global-wrds-basics/ provides various financial data of publicly traded companies in North America. Among them, we use the Annual Simplified Financial Statement, which includes key financial columns such as Company Name, Ticker Code, Year, Net Sales, and Total Assets, spanning multiple years for each company. The Annual Simplified Financial Statement consists of 80 columns, which are broadly classified into four categories: Identifying Information, Balance Sheet Variables, Income Statement Variables, and Statement of Cash Flows Variables. The components of each category are presented in Figure 24. This Annual Simplified Financial Statement will later be used to construct the table and chart modalities.
@@ B.2 Text Source
To construct texts that are closely related to the Annual Simplified Financial Statement of company entities, we focus on corporate financial reports. Companies listed on the U.S. stock market are required to periodically provide financial reports to the U.S. Securities and Exchange Commission (SEC), and these reports can be publicly accessed through the Electronic Data Gathering, Analysis, and Retrieval System (EDGAR).https://www.sec.gov/search-filings We use the annual disclosure report, the 10-K report, of companies to match the Annual Simplified Financial Statement. This 10-K report differs from the summary-style annual reports typically used in datasets such as FinQA Chen et al. (2021b) and TAT-QA Zhu et al. (2021), as it provides more in-depth financial data and disclosures. All companies’ 10-K reports are composed of a common table of contents format.
@@
Among the various items of the 10-K report, we use ITEM 1 (Business), ITEM 2 (Legal Proceedings), ITEM 7 (Management’s Discussion and Analysis of Financial Condition and Results of Operations), ITEM 7A (Quantitative and Qualitative Disclosures about Market Risk), and ITEM 8 (Financial Statements and Supplementary Data), which are most commonly used in actual corporate analysis. Further details can be found in the document provided by the SEChttps://www.sec.gov/files/reada10k.pdf.
@@ D.7 Problem Understanding & Condition Ignoring
Statement Types Example FC In May 2020, WORLD KINECT CORP modified and refreshed its asset-backed debt financing facility. CT The firm COSTCO WHOLESALE CORP. disclosed that its cogs values surpassed 92,765 during the year 2021. AR The company where the 2023 txt value minus the 2022 txt value equals 359.0 is WORLD KINECT CORP. TR Over the period from 2020 through 2023, SALESFORCE INC. consistently experienced an increase in its xsga values. RK During 2019, US FOODS HOLDING CORP. possessed the lowest txt value compared to other companies. FC+CT In 2023, the company whose xint values are less than 849.5 owns a terminal facility located at Fort Mifflin, which includes two docks for ships and has a total storage capacity of approximately 570 MBbls. FC+AR The company where the 2020 spi value minus the 2022 spi value equals 1036.0 recorded \$39 million and \$40 million in prior service credit amortization in 2018 and 2017, respectively. FC+TR The firm that showed a steady increase in dpc values between 2019 and 2023 must comply with the detailed regulations set by the Department of Transportation (DOT) regarding its pipeline infrastructure. FC+RK In 2022, the company, which reports fincf values greater than -2009, spreads out the amortization of its capitalized costs tied to new revenue contracts across four years. CT+TR Over the years, MONDELEZ INTERNATIONAL INC. consistently reports a fincf value exceeding -19575.5, while the ceq value demonstrates a continuous increase. CT+RK When the seq figure for WARNER BROS DISCOVERY INC falls below 10177.5, VERIZON COMMUNICATIONS INC records the lowest cogs value. AR+TR Throughout all periods, ENTERPRISE PRODCT PARTNRS LP is the company in which the cumulative nopi values surpass 1553.35, while the ao values have persistently increased. AR+RK During the years when lt value minus ibc value for UNITED PARCEL SERVICE INC is greater than 55265, the company with the lowest sale value is UNITED PARCEL SERVICE INC. FC+CT+TR Professional services are provided by the organization, which has the nopi value below 2796 for all years and the aoloch value that consistently declines for all years, to help clients with digital transformations using Salesforce solutions. FC+CT+RK The firm, with the highest act value during the years when ENERGY TRANSFER LP’s intan value dips below 8059, acquired a controlling interest in USAC through a \$250 million cash transaction. FC+AR+TR The organization, with the cumulative sum of nopi values below 2840 and continuously increasing act values for every year, has provided put rights to certain consolidated subsidiaries. These put rights are omitted from the contractual obligations table due to unpredictability in payment. FC+AR+RK The business, with the minimal aoloch value in the years when the ivao value minus the ao value for VERIZON COMMUNICATIONS INC exceeds 9312.5, provides expert services to support customers in executing digital transformations leveraging Salesforce solutions.

### v2

@@ 7 Conclusion
We introduce FCMR, a new benchmark designed to evaluate the cross-modal multi-hop reasoning ability of MLLMs. We evaluate the performance of state-of-the-art MLLMs, revealing that current models struggle with reasoning across different modalities. In future work, we plan to develop optimized methods to enhance performance based on the observations and analyses from this study.
@@
Reliance on Manual Analysis
Part of the analysis in this work is based on the manual inspection of in-house researchers. While this was inevitable to guarantee a high-quality, in-depth investigation, future work may concentrate more on automated analysis. We also emphasize that, despite its cost, manual analysis is worth investigating, as demonstrated in Section 5.2 and Section 6.2.
@@ Appendix A Details of Material Cross-Modal Multi-Hop Reasoning (Material-CMR)
The proposed CMRGen framework is easily adaptable to various domains, and as an example, we applied it to the field of Material Science to create the Material Cross-Modal Multi-Hop Reasoning (Material-CMR) dataset. We construct a Table Source containing material properties and a Text Source describing the materials through The Materials Project. The Materials Project [^12] https://next-gen.materialsproject.org/ is an initiative that accelerates materials discovery by providing researchers with computational data and tools to predict material properties, enabling more targeted and efficient experimental research.
@@ B.1 Table Source
Considering the practicality of cross-modal multi-hop reasoning, we utilize annual financial statements, an essential element in real corporate analysis, as the table data source. Wharton Research Data Services (WRDS) Compustat [^13] https://wrds-www.wharton.upenn.edu/pages/grid-items/compustat-global-wrds-basics/ provides various financial data of publicly traded companies in North America. Among them, we use the Annual Simplified Financial Statement, which includes key financial columns such as Company Name, Ticker Code, Year, Net Sales, and Total Assets, spanning multiple years for each company. The Annual Simplified Financial Statement consists of 80 columns, which are broadly classified into four categories: Identifying Information, Balance Sheet Variables, Income Statement Variables, and Statement of Cash Flows Variables. The components of each category are presented in Figure 28. This Annual Simplified Financial Statement will later be used to construct the table and chart modalities.
@@ B.2 Text Source
To construct texts that are closely related to the Annual Simplified Financial Statement of company entities, we focus on corporate financial reports. Companies listed on the U.S. stock market are required to periodically provide financial reports to the U.S. Securities and Exchange Commission (SEC), and these reports can be publicly accessed through the Electronic Data Gathering, Analysis, and Retrieval System (EDGAR). [^14] https://www.sec.gov/search-filings We use the annual disclosure report, the 10-K report, of companies to match the Annual Simplified Financial Statement. This 10-K report differs from the summary-style annual reports typically used in datasets such as FinQA Chen et al. (2021b) and TAT-QA Zhu et al. (2021), as it provides more in-depth financial data and disclosures. All companies’ 10-K reports are composed of a common table of contents format.
@@
Among the various items of the 10-K report, we use ITEM 1 (Business), ITEM 2 (Legal Proceedings), ITEM 7 (Management’s Discussion and Analysis of Financial Condition and Results of Operations), ITEM 7A (Quantitative and Qualitative Disclosures about Market Risk), and ITEM 8 (Financial Statements and Supplementary Data), which are most commonly used in actual corporate analysis. Further details can be found in the document provided by the SEC [^15] https://www.sec.gov/files/reada10k.pdf.
@@ D.1 Trend Assessment Error
Dataset: FCMR (Hard) Image? Accuracy
Random Selection - 12.28
GPT-4o 14.71
24.37
@@ D.7 Problem Understanding & Condition Ignoring
Appendix E Chart Interpretability
Model Accuracy (%)
Claude 3.5 Zero-Shot COT Baseline 32
Mulberry-LLaVA-8B Yao et al. (2024) 12
Virgo 72B Du et al. (2025) 19
Gemini 2.0 Flash Thinking Deepmind (2025) 39
o1 OpenAI (2024b) 43
Table 9: Performance of o1-like MLLMs on 100 randomly selected FCMR Hard-level samples. Due to the high cost of o1-like models, we limited our evaluation to these 100 samples.
Figure 9: Fine-grained Stage-based Analysis of Preliminary Task Optimization results applying Modality Integration, 4-Stage Reasoning, and Self-Refine.
@@ Appendix E Chart Interpretability
Statement Types Example
FC In May 2020, WORLD KINECT CORP modified and refreshed its asset-backed debt financing facility.
CT The firm COSTCO WHOLESALE CORP. disclosed that its cogs values surpassed 92,765 during the year 2021.
AR The company where the 2023 txt value minus the 2022 txt value equals 359.0 is WORLD KINECT CORP.
TR Over the period from 2020 through 2023, SALESFORCE INC. consistently experienced an increase in its xsga values.
RK During 2019, US FOODS HOLDING CORP. possessed the lowest txt value compared to other companies.
FC+CT In 2023, the company whose xint values are less than 849.5 owns a terminal facility located at Fort Mifflin, which includes two docks for ships and has a total storage capacity of approximately 570 MBbls.
FC+AR The company where the 2020 spi value minus the 2022 spi value equals 1036.0 recorded \$39 million and \$40 million in prior service credit amortization in 2018 and 2017, respectively.
FC+TR The firm that showed a steady increase in dpc values between 2019 and 2023 must comply with the detailed regulations set by the Department of Transportation (DOT) regarding its pipeline infrastructure.
FC+RK In 2022, the company, which reports fincf values greater than -2009, spreads out the amortization of its capitalized costs tied to new revenue contracts across four years.
CT+TR Over the years, MONDELEZ INTERNATIONAL INC. consistently reports a fincf value exceeding -19575.5, while the ceq value demonstrates a continuous increase.
CT+RK When the seq figure for WARNER BROS DISCOVERY INC falls below 10177.5, VERIZON COMMUNICATIONS INC records the lowest cogs value.
AR+TR Throughout all periods, ENTERPRISE PRODCT PARTNRS LP is the company in which the cumulative nopi values surpass 1553.35, while the ao values have persistently increased.
AR+RK During the years when lt value minus ibc value for UNITED PARCEL SERVICE INC is greater than 55265, the company with the lowest sale value is UNITED PARCEL SERVICE INC.
FC+CT+TR Professional services are provided by the organization, which has the nopi value below 2796 for all years and the aoloch value that consistently declines for all years, to help clients with digital transformations using Salesforce solutions.
FC+CT+RK The firm, with the highest act value during the years when ENERGY TRANSFER LP’s intan value dips below 8059, acquired a controlling interest in USAC through a \$250 million cash transaction.
FC+AR+TR The organization, with the cumulative sum of nopi values below 2840 and continuously increasing act values for every year, has provided put rights to certain consolidated subsidiaries. These put rights are omitted from the contractual obligations table due to unpredictability in payment.
FC+AR+RK The business, with the minimal aoloch value in the years when the ivao value minus the ao value for VERIZON COMMUNICATIONS INC exceeds 9312.5, provides expert services to support customers in executing digital transformations leveraging Salesforce solutions.
@@
Figure 25: Modality Integration prompt.
Figure 26: 4-Stage Reasoning Strategy prompt.
Figure 27: Self-Refine prompt.